# Towards Optimizing with Large Language Models

### v1

@@ 6-2 Analysis
Sensitivity to data samples. It’s worth considering that the aforementioned results may be influenced by the inherent randomness in the generation of test samples. Previous research has indicated that Large Language Models (LLMs) may demonstrate preferences for particular numbers, words, and symbols [2], which can introduce a level of bias in their responses. Given the high sensitivity of LLMs to the input prompt, the initial starting points and data provided can exert a significant influence on their outputs. In essence, the impact of instruction description and data initialization should be carefully considered when interpreting the results of LLM-based experiments to ensure a more accurate assessment of their performance.
VII Conclusion
@@
In this paper, we present our in-depth examination of assessing Large Language Models (LLMs) within the realm of optimization. We investigate LLMs’ performance across four optimization tasks that necessitate their comprehension of algorithmic instructions and their ability to generate new solutions based on previous solutions and their corresponding values.
@@
Interestingly, in cases where samples have fewer data points(variables), it’s common to observe a higher variance in solution quality. The notable pattern of uncertainty is characterized by an initial surge followed by a gradual decrease as data size increases and persistence across tasks and models.
In conclusion, our exploration of Large Language Models (LLMs) for optimization has unveiled a host of unresolved questions that beckon future research. Chief among these challenges is the quest to mitigate the sensitivity of LLMs to both data value and size, a paramount concern for enhancing their practical utility.
APPENDIX
Appendix A Instruction Prompts
Objective Function Prompt :
@@ Appendix A Instruction Prompts
Gradient-Descent Prompt :
@@
Short Answer: After calculation, the next update point is (ŷ$1_{new}$, ŷ$2_{new}$, ŷ$3_{new}$, …..) =
Grid-Search Prompt (Create Grid Points) :
@@
Grid-Search Prompt (Select) :
@@
Hill-Climbing Prompt (Generate Neighbors) :
@@
Hill-Climbing Prompt (Select) :
@@
Black-Box Prompt :
@@
Appendix B Examples
For tasks except Black Box Optimization, we use the whole dialogue as a prompt. Take the gradient-descent task as an example, the first prompt will be the Define Loss Function section, the second prompt will be the Define Loss Function section plus the First Iteration section and so on.
Gradient-Descent :
@@ Appendix B Examples
Short Answer: After calculation, the next update point is (ŷ$1_{new}$, ŷ$2_{new}$, ŷ$3_{new}$) = (1.81, 3.12, 3.41)
Grid-Search :
@@
List : [(2, 3, 3)]
Hill-Climbing :
@@
As for the Black-Box task, the first prompt will be the Loss Function section and for the subsequent iterations, the prompt will be the Loss Function section and the prompt with new past results.
Black-Box Optimization :

### v3

@@ 7. Analysis and Discussion
LLMs are sensitive to numerical values. It’s worth considering that the aforementioned results may be influenced by the inherent randomness in the generation of test samples. Previous research has indicated that LLMs may demonstrate preferences for particular numbers, words, and symbols (Renda et al., 2023), which can introduce a level of bias in their responses. Given the high sensitivity of LLMs to the input prompt, the initial starting points and data provided can exert a significant influence on their outputs. In essence, the impact of instruction description and data initialization should be carefully considered when interpreting the results of LLM-based experiments to ensure a more accurate assessment of their performance.
Self-consistency prompting improves stability. In the Gradient-Descent task, we employ self-consistency technique (Wang et al., 2022), where we conduct five repetitions for each iteration and select the solution that emerges most frequently. From Figure 6, we can see that GPT-4 performances increase largely, and the confidence interval for both the policy-metric and goal-metric narrows, indicating improved stability and reliability. Nonetheless, this approach does not yield favorable outcomes when applied to GPT-turbo-3.5. This suggests the need for further investigation within the realm of variance reduction.
Figure 6. The confidence intervals for both the policy and goal metrics of GPT-4 narrow, indicating improved stability. A negative policy metric with a high goal metric signifies significant outperformance of the ground truth model with six-dimensional instances.
8. Conclusion and Future Directions
@@
In this paper, we present our in-depth examination of assessing Large Language Models within the realm of optimization, where LLM progressively generates new solutions to optimize an objective function. We investigate LLMs’ performance across four optimization tasks that necessitate their comprehension of algorithmic instructions and their ability to generate new solutions based on previous solutions and their corresponding values.
@@
We also consolidate several crucial insights derived from our experimental results and subject them to analysis. We find that pretrained knowledge dominates the optimization capability of LLMs, while they also possess a richer solution space in small-scale problems. Furthermore, we elaborate on the potential of LLMs as hybrid optimizers. These insights and analyses unveil a host of unresolved questions that warrant further research.
Appendix A Prompt Templates
User Prompt:
@@ Appendix A Prompt Templates
Figure 7. Example prompt for getting objective function.
User Prompt:
@@
Short Answer: After calculation, the next update point is (ŷ$1_{new}$, ŷ$2_{new}$, ŷ$3_{new}$, …..) = …
Figure 8. Example prompt for Gradient-Descent.
User Prompt :
@@
What are the combinations? Your answer includes two parts an explanation with calculation and a list containing all the combinations.
@@
Figure 9. Example prompt for Grid-Search (Create Grid Points)
User Prompt :
@@
Figure 10. Example prompt for Grid-Search (Select)
User Prompt :
@@
Figure 11. Example prompt for Hill-Climbing Prompt (Generate Neighbors)
User Prompt :
@@
Figure 12. Example prompt for Hill-Climbing Prompt (Select)
User Prompt :
@@
Figure 13. Example prompt for Black-Box
User Prompt :
@@
Short Answer: After calculation, the next update point is (ŷ$1_{new}$, ŷ$2_{new}$, ŷ$3_{new}$) = (1.81, 3.12, 3.41)…
Figure 14. Iterative prompt for Gradient-Descent
User Prompt :
@@
List : [(2, 3, 3)]…
Figure 15. Iterative prompt for Grid-Search
User Prompt :
@@
Figure 16. Iterative prompt for Hill-Climbing
User Prompt :